what is the volume of a cylinder to the nearest tenth of a cubic inch with a radius of 7 inches a height of 12 inches

Answer by stanbon(57377)

what is the volume of a cylinder to the nearest tenth of a cubic inch with a radius of 7 inches a height of 12 inches
--------
Volume = (Area of base)*(height)
----
V = (pi)*7^2*12
---
V = (pi)*49*12
---
V = 588(pi) in^3 = 1847.3 in^3
===================
Cheers,
Stan H.
=================